OpenStudy (anonymous):

What is the value of x when -8x - (8x) = -48?

OpenStudy (anonymous):

You should try combining like terms first..

OpenStudy (akashdeepdeb):

-8x - (8x) = -48 -2 (8x) = -48 2(8x) = 48 x = 48/(2*8) = 48/16

OpenStudy (anonymous):

-8x -8x is just like -1 -1 = -2..

OpenStudy (anonymous):

the answer was like 3 -.-